OpenStudy (anonymous):

You have a 4 gallon jug and a 7 gallon jug, how do you measure exactly 2 gallons of water? Someone please help!

OpenStudy (briannaontimeforreal):

Empty three gallons from the five gallon jug into the three gallon jug. ... MATH. Thus we have exactly one gallon. Also, we can obtain $11-3=8$ gallons. ... theory to a practical problem of capacity and achieving a certain volumetric measure. ... In this problem, you cannot try to solve for something directly; you must instead

OpenStudy (anonymous):

Theres not a 5 gallon jug :L

OpenStudy (briannaontimeforreal):

i meant to say 4

OpenStudy (campbell_st):

fill the 7 gallon... then use the 7 gallon to fill the 4 gallon you will have 3 gallons left in the 7 gallon next step empty the 4 gallon, and then pour in the 3 gallons remaining in the 7 gallon. next step refill the 7 gallon, use this to fill the 4 gallon, you will be ponly able to put in 1 gallon. leaving 6 gallons in the 7 gallon jug. Last step.... you should be able to work it out from here

OpenStudy (dayakar):

fill with 7 G jug 2 times , remove 3 times with 4 G jug = 2* 7 - 3* 4 = 14- 12

OpenStudy (anonymous):

How do I know to fill the 7 gallon twice and the 4 gallon 4 times.. ?

OpenStudy (anonymous):

@campbell_st I came that far aswell, I still don't understand how to finish it.

OpenStudy (campbell_st):

well you have to empty the 4 gallon... so you have an empty 4 gallon and a 7 gallon which contains 6 gallons... how do you get 2 gallons...? its not hard

OpenStudy (anonymous):

why not just fill the 4 gallon jug up half way?